The dipole moment of hydrogen chloride is 1.03 D. Its percent ionic character is about 17%. Its bond distance in pm is about?
 

Grade:12th pass

1 Answers

Pooja Kumari
7 years ago
Percentage ionic character= ( observed dipole moment / calculated dipole moment )*100 ; Calculated dipole moment = q * d ; Applying these 2 equation : calc. Dipole moment ~ 6; d = 6/ 4.8*(10)^(-10) ; d = 1.25 * (10)^(-10) cm